# Generative and Protective AI for Content Creation

### v1

@@ 1 Introduction
Generative audio models have undergone rapid advances in recent years, driven largely by diffusion (DMs) tango; tango2; audioldm; audioldm2 and autoregressive architectures (ARMs) figaro; audiogen. These models are capable of producing highly realistic soundscapes, speech, and music. While this progress opens exciting opportunities in areas such as creative expression, accessibility, and interactive media, it also raises urgent concerns about privacy, copyright, and data governance. In particular, the vast datasets required to train such systems can contain artistic or commercial audio without transparent disclosure, leaving creators uncertain about whether their works contributed to a model’s capabilities dornis2025generative. A central question in this context is whether one can reliably determine if a specific artist’s recordings were included in training of a generative model. Addressing this question is critical both for protecting intellectual property and for enabling accountability in machine learning practice. Similar challenges have been investigated in computer vision kong2023efficient; dubinski2024towards; carlini2023extracting and natural language processing maini2024llmdatasetinferencedid; mattern2023membershipLLM; shi2024detecting, where MIAs shokri2017membershipinference attempt to determine if a given sample was used in training, and DI maini2021dataset; maini2024llmdatasetinferencedid; dubinski2025cdi extends this idea to entire collections. However, the effectiveness of these techniques for large generative models in the audio domain remains unclear.
@@ 2.1 Identifying Training Data
Membership Inference (MIA). MIAs (shokri2017membershipinference) aim to decide whether a given sample was part of a model’s training set. They exploit overfitting: training samples typically yield lower losses than unseen ones. Formally, the attacker constructs an attack function $A_{f_{\theta}}:\mathcal{X}\rightarrow{0,1}$ that predicts membership. A standard approach is the threshold attack (yeom2018lossmia), which classifies $x$ as a member if the chosen metric is below a threshold: $A_{f_{\theta}}(x)=\mathbb{1}!\left[\mathcal{M}(f_{\theta},x)<\gamma\right]$, where $\mathcal{M}$ is the metric and $\gamma$ the decision threshold.
@@
Dataset Inference (DI). DI (maini2021dataset) asks whether an entire dataset was used during training. Unlike MIAs, which evaluate individual samples, DI aggregates membership signals (often based on MIAs) across multiple points into a dataset-level statistic, thereby amplifying weak per-sample evidence. This makes DI effective for large models and datasets where single-sample inference is unreliable. Initially proposed for supervised models, DI extracts per-sample features, aggregates them into a dataset score, and applies a statistical test (sslextractions2022icml; datasetinference2022neurips). Recent work has extended DI to generative models, including large language models (LLMs) (maini2024llmdatasetinferencedid; zhao2025posthocDI), DMs (dubinski2025cdi), and autoregressive image models (kowalczukprivacy). Formally, DI compares scores from a suspected member set and a non-member set via Welch’s $t$-test at $\alpha=0.01$ with the null hypothesis $H_{0}:\text{mean(member scores)}\leq\text{mean(non-member scores)}$. Rejecting $H_{0}$ implies the dataset was part of training. Correctness requires both sets be i.i.d.; otherwise, distributional mismatch can bias the test. The strength of DI depends on the number of available samples. To quantify leakage, we define $P$ as the minimum number of samples needed to reject $H_{0}$. Smaller $P$ indicates stronger leakage.
@@
Figure 1: Dataset Inference Procedural Steps. The process consists of four main steps: 1 Data Preparation: Prepare the data to verify whether the (suspected) member samples P were used to train the model. The (confirmed) nonmember samples U, from the same distribution as P, serve as the validation set. 2 Feature Extraction: Run each individual MIA on all inputs from $\{\textbf{P},\textbf{U}\}$ to extract membership features for all data samples. We use our MIAs tailored to ARMs and DMs for the respective model types. 3 Score Computation: We map extracted feature vectors into scalar membership scores for each sample. We use a scoring model for DMs, following dubinski2025cdi, and feature summation kowalczukprivacy for ARMs (see Appendix for more details. 4 Statistical Testing: Apply a statistical t-test to verify whether the scores obtained for the public suspect data points P are statistically significantly higher than those for U. If so, P is marked as being part of the model’s training set. Otherwise, the test is inconclusive and the model’s training set is considered independent of P.
@@ 2.2 Audio Generative Models
We experiment on 4 models described in Table˜1 representing both AR and DM famielies. Currently, DMs dominate high-quality audio generation. AudioLDM2 audioldm2 unifies text-to-audio, text-to-music, and text-to-speech within one framework. It leverages a "Language of Audio" representation, mapped from AudioMAE audiomae_huang2022amae features through GPT-2 radford2019language, to condition a UNet ronneberger2015unet diffusion model over mel-spectrogram latents. With 29.5k hours of diverse training data, it establishes a common semantic space that supports multiple generative tasks. TANGO (tango) follows a simpler design. A frozen FLAN-T5 chung2022scaling encoder provides text embeddings that guide a latent diffusion model trained on the AudioCaps audiocaps dataset. Augmentation is based on pressure-level mixing of sounds, ensuring balanced exposure. Despite training on only 46k clips, TANGO achieves competitive generation quality by relying on strong instruction-tuned text features.
Competing with DMs across different modalities, autoregressive approaches have also been applied to audio generation. AudioGen audiogen treats audio as a sequence of discrete codec tokens and learns a text-conditioned Transformer decoder to generate them. The model is trained on roughly 4k hours from ten heterogeneous audio–text corpora, and augmentations based on mixing sound sources are used to expose it to overlapping events. FIGARO figaro addresses symbolic music with controllable generation. It introduces description-to-sequence learning, combining interpretable features such as chords, instrumentation, and rhythm with learned latent codes as conditioning. Training on the 176k-file LakhMIDI dataset raffel2016learning allows the model to reconstruct bar-level sequences and to provide global and fine-grained control.
@@ 3 Method
Membership Inference. We begin by evaluating MIAs on large audio generative models. For DMs, we apply the attack suite explored in dubinski2025cdi, which exploits denoising dynamics to distinguish training samples from non-members. For ARMs, we use the approach from Kowalczuk et al. kowalczukprivacy, which leverages token-level log-likelihoods and related statistics. We give more details on individual MIAs used in our work in the Appendix. In both cases, we use the train split of each model’s dataset as the source of members and the held-out test split as the source of non-members. This ensures that attacks are evaluated under a realistic and controlled setting where the attacker’s candidate pool contains both genuine training data and independent test samples.
Dataset Inference. To facilitate the task, we extend our studies to DI. We follow the methodology introduced by dubinski2025cdi for DMs and kowalczukprivacy for ARMs. Each candidate dataset consists of a collection of suspected member samples $\mathcal{P}$ and an equal number of non-member samples $\mathcal{U}$ drawn from the test split. We extract membership features for each sample using multiple MIAs, aggregate them into scalar scores, and then apply Welch’s $t$-test to compare $\mathcal{P}$ and $\mathcal{U}$. The null hypothesis states that the mean score of $\mathcal{P}$ is no greater than that of $\mathcal{U}$, and we reject it at $\alpha=0.01$ if sufficient evidence is found. Following standard practice, we report the minimum number of samples $P$ required to reject the null hypothesis, with smaller $P$ indicating stronger information leakage. Our approach is demonstrated on Figure˜1.
@@ Acknowledgments
This work is partly supported by the MUR FIS2 grant n. FIS-2023-00942 "NEXUS" (cup B53C25001030001), and by Sapienza University of Rome via the Seed of ERC grant "MINT.AI" (cup B83C25001040001).
@@ Appendix A Details on Membership Inference for Autoregressive Models
Kowalczuk et al. [kowalczukprivacy] introduce the first comprehensive MIA suite for image autoregressive models by adapting well-established, token-level attacks from the LLM literature (e.g., Loss, Zlib, Hinge, Min-K%, Min-K%++, SURP, CAMIA) to visual next-token prediction. A key observation in their work is that many IARs are trained with classifier-free guidance [ho2022classifier], i.e., the forward pass processes each example both with conditioning (e.g., a class label or text prompt) and without it. Building on CLiD [zhai2024clid], they exploit this extra supervision signal by contrasting the conditional and unconditional paths: instead of feeding raw per-token logits into MIAs, they use the guidance-difference statistic
@@
Min-k% Prob metric. To reduce the influence of highly predictable positions, [shi2024detecting] focus the decision rule on the least likely part of the sequence. For an input $x$ and a fraction $K\!\in\!\{10,20,30,40,50\}$, Min-k% Prob computes the average negative log-likelihood over the bottom $K\%$ tokens under the model $f_{\theta}$. Membership is predicted by thresholding this average:
@@
Zlib ratio attack. This baseline relates model fit to a model-agnostic compressibility proxy [zlib2004]. Let $\mathcal{P}_{f_{\theta}}(x)$ denote the perplexity (or exponentiated average negative log-likelihood) and $\operatorname{zlib}(x)$ be the compressed size of $x$ under the zlib codec. The statistic
@@
CAMIA. Context-aware MIA [chang2024context] augments raw loss features with temporal descriptors of the token-wise loss sequence. Several signals are used: a slope feature that captures how quickly losses decline across positions; approximate entropy, which measures regularity by the prevalence of repeating patterns; Lempel–Ziv complexity, which quantifies diversity in the loss trajectory via the count of distinct substrings; a count-below statistic, the fraction of tokens with loss below a preset cutoff; and a repeated-sequence amplification feature that measures the reduction in loss when the same input is repeated. Non-members typically display higher irregularity and larger gains from repetition, while members show more stable, low-loss segments.
@@ Appendix B Details on Membership Inference for Diffusion Models
Diffusion Models [songdenoising] are generative models trained by progressively adding noise to the data and then learning to reverse this corruption. In the forward diffusion process, Gaussian noise $\epsilon\sim\mathcal{N}(0,I)$ is added to a clean sample $x$ to obtain a noised sample $x_{t}\leftarrow\sqrt{\alpha_{t}}\,x+\sqrt{1-\alpha_{t}}\,\epsilon$, where $t\in[0,T]$ is the diffusion timestep and $\alpha_{t}\in[0,1]$ is a monotonically decreasing schedule with $\alpha_{0}=1$ and $\alpha_{T}=0$. The denoiser $f_{\theta}$ is trained to predict $\epsilon$ across timesteps by minimizing $\frac{1}{N}\sum_{i}\mathbb{E}_{t,\epsilon}\,\mathcal{L}(x_{i},t,\epsilon;f_{\theta})$, where $N$ is the training set size and
@@
Latent diffusion models [rombach2022high] perform the diffusion process in a learned latent space to improve efficiency. An encoder $\mathcal{E}$ maps $x$ to a latent $z=\mathcal{E}(x)$, and the objective in Eq. 9 becomes
@@
Denoising Loss. Early membership inference attacks for diffusion models [carlini2023extracting] assess sample membership by directly using the denoising loss as a statistic. The key observation is that the loss at intermediate timesteps provides the strongest separation between training members and non-members. In particular, $t\approx 100$ often yields the most discriminative signal: very small $t$ makes the task too easy (the noised input remains close to the original), whereas very large $t$ collapses the input toward pure noise, making prediction uniformly hard. A sample is classified as a member if its loss at the chosen timestep falls below a threshold selected on a validation split.
@@
Proximal Initialization Attack (PIA). The PIA family [kong2023efficient] compares the model’s noise predictions when initialized from different proximity states to the data. A canonical instantiation evaluates the prediction error at a clean (or minimally noised) state, such as $t=0$, and at a moderately noised state, typically around $t=200$ where separability is reported to be strong. The difference (or ratio) between these errors is used as the attack feature. Intuitively, training samples induce more confident and stable predictions across nearby states of the diffusion process, leading to a lower feature value for members than for non-members.
PIAN. An adaptation of PIA, denoted PIAN [kong2023efficient], normalizes the denoiser’s output to enforce approximately Gaussian behavior in the predicted noise, thereby reducing scale effects that may confound raw error magnitudes. The membership statistic is computed analogously to PIA after normalization. As with PIA, members are expected to yield smaller scores because the model’s predictions align more consistently with the true noise on training data.
@@
The gradient-masking attack [dubinski2025cdi] targets semantically critical regions of the latent representation that most influence the denoising loss. For a given $z_{t}$, the gradient $\mathbf{g}=\left\lvert\nabla_{z_{t}}\,\mathcal{L}(z_{t},t,\epsilon;f_{\theta})\right\rvert$ is computed, and a binary mask $\mathbf{M}$ is formed by selecting the top-percentile (e.g., top $20\%$) entries of $\mathbf{g}$. A perturbed latent $\hat{z}_{t}=\epsilon\cdot\mathbf{M}+z_{t}\cdot\neg\mathbf{M}$ is then created by replacing the most influential coordinates with random noise and leaving the remainder unchanged. The attack feature is the reconstruction error restricted to the masked region, $\left\lVert(\epsilon-z_{t})\cdot\mathbf{M}-f_{\theta}(\hat{z}_{t},t)\cdot\mathbf{M}\right\rVert_{2}^{2}$, optionally aggregated across multiple timesteps. Because models tend to memorize salient structure in training samples, members exhibit lower masked-region reconstruction error than non-members.
@@
The central premise of noise-optimisation attack [dubinski2025cdi] is that stronger (or more effective) perturbations are required to significantly reduce the denoising loss for training members, reflecting higher confidence and tighter fit on seen data. Concretely, starting from $z_{t}$ at an intermediate timestep (e.g., $t=100$), an unconstrained optimization seeks a perturbation $\delta$ that minimizes the objective $\min_{\delta}\,\left\lVert\epsilon-f_{\theta}(z_{t}+\delta,t)\right\rVert_{2}^{2}$, using 5 L-BFGS steps. Two complementary features arise: the minimized prediction error $\left\lVert\epsilon-f_{\theta}(z_{t}+\delta,t)\right\rVert_{2}^{2}$ and the perturbation magnitude $\lVert\delta\rVert_{2}^{2}$. Members typically achieve lower final error yet require larger or more targeted adjustments, producing distinctive signatures relative to non-members.

### v2

@@ 1 Introduction
Generative audio models have undergone rapid advances in recent years, driven largely by diffusion (DMs) Ghosal et al. (2023); Majumder et al. (2024); Liu et al. (2023, 2024) and autoregressive architectures (ARMs) von Rütte et al. (2023); Kreuk et al. (2022). These models are capable of producing highly realistic soundscapes, speech, and music. While this progress opens exciting opportunities in areas such as creative expression, accessibility, and interactive media, it also raises urgent concerns about privacy, copyright, and data governance. In particular, the vast datasets required to train such systems can contain artistic or commercial audio without transparent disclosure, leaving creators uncertain about whether their works contributed to a model’s capabilities Dornis and Stober (2025). A central question in this context is whether one can reliably determine if a specific artist’s recordings were included in training of a generative model. Addressing this question is critical both for protecting intellectual property and for enabling accountability in machine learning practice. Similar challenges have been investigated in computer vision Kong et al. (2023); Dubiński et al. (2024); Carlini et al. (2023) and natural language processing Maini et al. (2024); Mattern et al. (2023); Shi et al. (2024), where MIAs Shokri et al. (2017) attempt to determine if a given sample was used in training, and DI Maini et al. (2021, 2024); Dubiński et al. (2025) extends this idea to entire collections. However, the effectiveness of these techniques for large generative models in the audio domain remains unclear.
@@ 2.1 Identifying Training Data
Membership Inference (MIA). MIAs (Shokri et al., 2017) aim to decide whether a given sample was part of a model’s training set. They exploit overfitting: training samples typically yield lower losses than unseen ones. Formally, the attacker constructs an attack function $A_{f_{\theta}}:\mathcal{X}\rightarrow{0,1}$ that predicts membership. A standard approach is the threshold attack (Yeom et al., 2018), which classifies $x$ as a member if the chosen metric is below a threshold: $A_{f_{\theta}}(x)=\mathbb{1}!\left[\mathcal{M}(f_{\theta},x)<\gamma\right]$, where $\mathcal{M}$ is the metric and $\gamma$ the decision threshold.
@@
Dataset Inference (DI). DI (Maini et al., 2021) asks whether an entire dataset was used during training. Unlike MIAs, which evaluate individual samples, DI aggregates membership signals (often based on MIAs) across multiple points into a dataset-level statistic, thereby amplifying weak per-sample evidence. This makes DI effective for large models and datasets where single-sample inference is unreliable. Initially proposed for supervised models, DI extracts per-sample features, aggregates them into a dataset score, and applies a statistical test (Dziedzic et al., 2022a, b). Recent work has extended DI to generative models, including large language models (LLMs) (Maini et al., 2024; Zhao et al., 2025), DMs (Dubiński et al., 2025), and autoregressive image models (Kowalczuk et al., ). Formally, DI compares scores from a suspected member set and a non-member set via Welch’s $t$-test at $\alpha=0.01$ with the null hypothesis $H_{0}:\text{mean(member scores)}\leq\text{mean(non-member scores)}$. Rejecting $H_{0}$ implies the dataset was part of training. Correctness requires both sets be i.i.d.; otherwise, distributional mismatch can bias the test. The strength of DI depends on the number of available samples. To quantify leakage, we define $P$ as the minimum number of samples needed to reject $H_{0}$. Smaller $P$ indicates stronger leakage.
@@
Figure 1: Dataset Inference Procedural Steps. The process consists of four main steps: 1 Data Preparation: Prepare the data to verify whether the (suspected) member samples P were used to train the model. The (confirmed) nonmember samples U, from the same distribution as P, serve as the validation set. 2 Feature Extraction: Run each individual MIA on all inputs from $\{\textbf{P},\textbf{U}\}$ to extract membership features for all data samples. We use our MIAs tailored to ARMs and DMs for the respective model types. 3 Score Computation: We map extracted feature vectors into scalar membership scores for each sample. We use a scoring model for DMs, following Dubiński et al. (2025), and feature summation Kowalczuk et al. for ARMs (see Appendix for more details. 4 Statistical Testing: Apply a statistical t-test to verify whether the scores obtained for the public suspect data points P are statistically significantly higher than those for U. If so, P is marked as being part of the model’s training set. Otherwise, the test is inconclusive and the model’s training set is considered independent of P.
@@ 2.2 Audio Generative Models
We experiment on 4 models described in Table˜1 representing both AR and DM famielies. Currently, DMs dominate high-quality audio generation. AudioLDM2 Liu et al. (2024) unifies text-to-audio, text-to-music, and text-to-speech within one framework. It leverages a "Language of Audio" representation, mapped from AudioMAE Huang et al. (2022) features through GPT-2 Radford et al. , to condition a UNet Ronneberger et al. (2015) diffusion model over mel-spectrogram latents. With 29.5k hours of diverse training data, it establishes a common semantic space that supports multiple generative tasks. TANGO (Ghosal et al., 2023) follows a simpler design. A frozen FLAN-T5 Chung et al. (2022) encoder provides text embeddings that guide a latent diffusion model trained on the AudioCaps Kim et al. (2019) dataset. Augmentation is based on pressure-level mixing of sounds, ensuring balanced exposure. Despite training on only 46k clips, TANGO achieves competitive generation quality by relying on strong instruction-tuned text features.
Competing with DMs across different modalities, autoregressive approaches have also been applied to audio generation. AudioGen Kreuk et al. (2022) treats audio as a sequence of discrete codec tokens and learns a text-conditioned Transformer decoder to generate them. The model is trained on roughly 4k hours from ten heterogeneous audio–text corpora, and augmentations based on mixing sound sources are used to expose it to overlapping events. FIGARO von Rütte et al. (2023) addresses symbolic music with controllable generation. It introduces description-to-sequence learning, combining interpretable features such as chords, instrumentation, and rhythm with learned latent codes as conditioning. Training on the 176k-file LakhMIDI dataset Raffel (2016) allows the model to reconstruct bar-level sequences and to provide global and fine-grained control.
@@ 3 Method
Membership Inference. We begin by evaluating MIAs on large audio generative models. For DMs, we apply the attack suite explored in Dubiński et al. (2025), which exploits denoising dynamics to distinguish training samples from non-members. For ARMs, we use the approach from Kowalczuk et al. Kowalczuk et al. , which leverages token-level log-likelihoods and related statistics. We give more details on individual MIAs used in our work in the Appendix. In both cases, we use the train split of each model’s dataset as the source of members and the held-out test split as the source of non-members. This ensures that attacks are evaluated under a realistic and controlled setting where the attacker’s candidate pool contains both genuine training data and independent test samples.
Dataset Inference. To facilitate the task, we extend our studies to DI. We follow the methodology introduced by Dubiński et al. (2025) for DMs and Kowalczuk et al. for ARMs. Each candidate dataset consists of a collection of suspected member samples $\mathcal{P}$ and an equal number of non-member samples $\mathcal{U}$ drawn from the test split. We extract membership features for each sample using multiple MIAs, aggregate them into scalar scores, and then apply Welch’s $t$-test to compare $\mathcal{P}$ and $\mathcal{U}$. The null hypothesis states that the mean score of $\mathcal{P}$ is no greater than that of $\mathcal{U}$, and we reject it at $\alpha=0.01$ if sufficient evidence is found. Following standard practice, we report the minimum number of samples $P$ required to reject the null hypothesis, with smaller $P$ indicating stronger information leakage. Our approach is demonstrated on Figure˜1.
@@ Acknowledgments
This research was supported by the Polish National Science Centre (NCN) within grant no. 2023/51/I/ST6/02854.We gratefully acknowledge Poland’s high-performance Infrastructure PLGrid for providing computer facilities and support within computational grant no. PLG/2025/018230. This work is also partly supported by the MUR FIS2 grant n. FIS-2023-00942 "NEXUS" (cup B53C25001030001), and by Sapienza University of Rome via the Seed of ERC grant "MINT.AI" (cup B83C25001040001). This work was also supported by the German Research Foundation (DFG) within the framework of the Weave Programme under the project titled "Protecting Creativity: On the Way to Safe Generative Models" with number 545047250.
@@ Appendix A Details on Membership Inference for Autoregressive Models
Kowalczuk et al. [17] introduce the first comprehensive MIA suite for image autoregressive models by adapting well-established, token-level attacks from the LLM literature (e.g., Loss, Zlib, Hinge, Min-K%, Min-K%++, SURP, CAMIA) to visual next-token prediction. A key observation in their work is that many IARs are trained with classifier-free guidance [13], i.e., the forward pass processes each example both with conditioning (e.g., a class label or text prompt) and without it. Building on CLiD [34], they exploit this extra supervision signal by contrasting the conditional and unconditional paths: instead of feeding raw per-token logits into MIAs, they use the guidance-difference statistic
@@
Min-k% Prob metric. To reduce the influence of highly predictable positions, [29] focus the decision rule on the least likely part of the sequence. For an input $x$ and a fraction $K\!\in\!\{10,20,30,40,50\}$, Min-k% Prob computes the average negative log-likelihood over the bottom $K\%$ tokens under the model $f_{\theta}$. Membership is predicted by thresholding this average:
@@
Zlib ratio attack. This baseline relates model fit to a model-agnostic compressibility proxy [11]. Let $\mathcal{P}_{f_{\theta}}(x)$ denote the perplexity (or exponentiated average negative log-likelihood) and $\operatorname{zlib}(x)$ be the compressed size of $x$ under the zlib codec. The statistic
@@
CAMIA. Context-aware MIA [4] augments raw loss features with temporal descriptors of the token-wise loss sequence. Several signals are used: a slope feature that captures how quickly losses decline across positions; approximate entropy, which measures regularity by the prevalence of repeating patterns; Lempel–Ziv complexity, which quantifies diversity in the loss trajectory via the count of distinct substrings; a count-below statistic, the fraction of tokens with loss below a preset cutoff; and a repeated-sequence amplification feature that measures the reduction in loss when the same input is repeated. Non-members typically display higher irregularity and larger gains from repetition, while members show more stable, low-loss segments.
@@ Appendix B Details on Membership Inference for Diffusion Models
Diffusion Models [31] are generative models trained by progressively adding noise to the data and then learning to reverse this corruption. In the forward diffusion process, Gaussian noise $\epsilon\sim\mathcal{N}(0,I)$ is added to a clean sample $x$ to obtain a noised sample $x_{t}\leftarrow\sqrt{\alpha_{t}}\,x+\sqrt{1-\alpha_{t}}\,\epsilon$, where $t\in[0,T]$ is the diffusion timestep and $\alpha_{t}\in[0,1]$ is a monotonically decreasing schedule with $\alpha_{0}=1$ and $\alpha_{T}=0$. The denoiser $f_{\theta}$ is trained to predict $\epsilon$ across timesteps by minimizing $\frac{1}{N}\sum_{i}\mathbb{E}_{t,\epsilon}\,\mathcal{L}(x_{i},t,\epsilon;f_{\theta})$, where $N$ is the training set size and
@@
Latent diffusion models [27] perform the diffusion process in a learned latent space to improve efficiency. An encoder $\mathcal{E}$ maps $x$ to a latent $z=\mathcal{E}(x)$, and the objective in Eq. 9 becomes
@@
Denoising Loss. Early membership inference attacks for diffusion models [3] assess sample membership by directly using the denoising loss as a statistic. The key observation is that the loss at intermediate timesteps provides the strongest separation between training members and non-members. In particular, $t\approx 100$ often yields the most discriminative signal: very small $t$ makes the task too easy (the noised input remains close to the original), whereas very large $t$ collapses the input toward pure noise, making prediction uniformly hard. A sample is classified as a member if its loss at the chosen timestep falls below a threshold selected on a validation split.
@@
Proximal Initialization Attack (PIA). The PIA family [16] compares the model’s noise predictions when initialized from different proximity states to the data. A canonical instantiation evaluates the prediction error at a clean (or minimally noised) state, such as $t=0$, and at a moderately noised state, typically around $t=200$ where separability is reported to be strong. The difference (or ratio) between these errors is used as the attack feature. Intuitively, training samples induce more confident and stable predictions across nearby states of the diffusion process, leading to a lower feature value for members than for non-members.
PIAN. An adaptation of PIA, denoted PIAN [16], normalizes the denoiser’s output to enforce approximately Gaussian behavior in the predicted noise, thereby reducing scale effects that may confound raw error magnitudes. The membership statistic is computed analogously to PIA after normalization. As with PIA, members are expected to yield smaller scores because the model’s predictions align more consistently with the true noise on training data.
@@
The gradient-masking attack [7] targets semantically critical regions of the latent representation that most influence the denoising loss. For a given $z_{t}$, the gradient $\mathbf{g}=\left\lvert\nabla_{z_{t}}\,\mathcal{L}(z_{t},t,\epsilon;f_{\theta})\right\rvert$ is computed, and a binary mask $\mathbf{M}$ is formed by selecting the top-percentile (e.g., top $20\%$) entries of $\mathbf{g}$. A perturbed latent $\hat{z}_{t}=\epsilon\cdot\mathbf{M}+z_{t}\cdot\neg\mathbf{M}$ is then created by replacing the most influential coordinates with random noise and leaving the remainder unchanged. The attack feature is the reconstruction error restricted to the masked region, $\left\lVert(\epsilon-z_{t})\cdot\mathbf{M}-f_{\theta}(\hat{z}_{t},t)\cdot\mathbf{M}\right\rVert_{2}^{2}$, optionally aggregated across multiple timesteps. Because models tend to memorize salient structure in training samples, members exhibit lower masked-region reconstruction error than non-members.
@@
The central premise of noise-optimisation attack [7] is that stronger (or more effective) perturbations are required to significantly reduce the denoising loss for training members, reflecting higher confidence and tighter fit on seen data. Concretely, starting from $z_{t}$ at an intermediate timestep (e.g., $t=100$), an unconstrained optimization seeks a perturbation $\delta$ that minimizes the objective $\min_{\delta}\,\left\lVert\epsilon-f_{\theta}(z_{t}+\delta,t)\right\rVert_{2}^{2}$, using 5 L-BFGS steps. Two complementary features arise: the minimized prediction error $\left\lVert\epsilon-f_{\theta}(z_{t}+\delta,t)\right\rVert_{2}^{2}$ and the perturbation magnitude $\lVert\delta\rVert_{2}^{2}$. Members typically achieve lower final error yet require larger or more targeted adjustments, producing distinctive signatures relative to non-members.